OpenStudy (anonymous):

find the tangential and normal components of the acceleration vector: r(t) = (cosT)i +(sinT)j + (T)K

OpenStudy (amistre64):

looks familiar

OpenStudy (amistre64):

the tangent and normal are unit lngths tho right?

OpenStudy (amistre64):

r' = tangent r'' = normal

OpenStudy (anonymous):

ummm i know the formula for tangent is (v dot a)/magnitude v

OpenStudy (amistre64):

that just gives you a scalar, not a vector

OpenStudy (amistre64):

v.a is a number and |v| is a number; so they dont produce a vector

OpenStudy (anonymous):

the answer is 0 for tangential and 1 for normal, i just don't know how they got it

OpenStudy (amistre64):

can you type up the generic form of an acceleration vector? I know its got something like dT/dt + kNdT/dt or some such

OpenStudy (anonymous):

umm i think its a=v'T = k(v^2)N?

OpenStudy (amistre64):

where T and N are units tan and norm; gonna have to check the google on that tho :)

OpenStudy (anonymous):

i meant a + in-between, not =

OpenStudy (anonymous):

you use this formula to solve it: but i just can't get the right answer, can you see what you get when you plug it in

OpenStudy (amistre64):

\[a_T=\frac{r'.r''}{|r'|}\] \[a_N=\frac{|r'xr''|}{|r'|}\]

OpenStudy (anonymous):

\[a _{T} = r'(T) * r''(T)/\left|r'(t)\right|\]

OpenStudy (anonymous):

yes with that formula

OpenStudy (anonymous):

and you're supposed to get 0 and 1 but I'm not getting either of those

OpenStudy (amistre64):

r(t) = < cosT, sinT, T > r'(t) = < -sinT, cosT, 1 > r''(t) = < -cosT, -sinT, 0 > we agree on that?

OpenStudy (anonymous):

yup

OpenStudy (amistre64):

r'(t) = < -sinT, cosT, 1 > .r''(t) = < -cosT, -sinT, 0 > ------------------------- sincos - sincos + 0 = 0 so At = 0 regardless of |r'|

OpenStudy (anonymous):

ohhhhh ohhhh i had plus 1k instead of 0

OpenStudy (amistre64):

r'(t) = < -sinT, cosT, 1 > x r''(t) = < -cosT, -sinT, 0 > -------------------------- 0 + sinT = sinT -(0+cos(t)) = -cosT sin^2 + cos^2 = 1 we agree with the cross?

OpenStudy (amistre64):

since the cross is the same components as r'; its |cross|/|r'| = 1

OpenStudy (anonymous):

yeah

OpenStudy (amistre64):

sin^2+cos^2+1^2 ---------------- = 1 sin^2+cos^2+1^2

OpenStudy (anonymous):

you got 1 for for magnitude r'(t)?

OpenStudy (amistre64):

doesnt matter what |r'| actually is; since its larger than 0

OpenStudy (amistre64):

we have the same results top to bottom so anyting/itself = 1

OpenStudy (anonymous):

i got |r'(t)| = 2 though

OpenStudy (anonymous):

actually i got root2

OpenStudy (amistre64):

then |r'xr''| will be 2 and 2/2 = 1

OpenStudy (amistre64):

in this instance: |r'xr''| = |r'| = a a/a = 1

OpenStudy (anonymous):

why does |r'xr''| = |r'|

OpenStudy (amistre64):

what are your compenets for r'xr'' ?

OpenStudy (amistre64):

components ..

OpenStudy (anonymous):

(-sin t, cos t, 1) x (-cos t, -sin t, 0)

OpenStudy (amistre64):

thats not the end result; that is what its made from

OpenStudy (anonymous):

hang on let me do it really quick

OpenStudy (amistre64):

when you actually cross them, i did this above, what do we get as components?

OpenStudy (anonymous):

sin, -cos, sin^2 + cos^2

OpenStudy (amistre64):

good, and s^2+c^2 = 1 right?

OpenStudy (anonymous):

yup

OpenStudy (amistre64):

now, look at the components of r' and tell me if there is any difference

OpenStudy (anonymous):

just the negative signs

OpenStudy (anonymous):

but you square it!

OpenStudy (anonymous):

so it won't matter!

OpenStudy (amistre64):

cross = sin, -cos, 1 r' = sin, cos, 1 when we take the magnitude of each we square the components, add them and sqrt them cross r' sin^2 = sin^2 ..... correct

OpenStudy (anonymous):

ahhhh genius! thank you so much!

OpenStudy (amistre64):

youre welcome :)